OpenStudy (anonymous):

ok guys i need help.... how do you do verify this trig identity? cscx (cscx-sinx) + (sinx-cosx)/sinx + cotx = csc squared x

myininaya (myininaya):

split that 2nd fraction into -cotx+cotx=0

OpenStudy (anonymous):

ok...

myininaya (myininaya):

i mean do you see that 2nd fraction beomces 1-cotx and after that we see +cotx

OpenStudy (anonymous):

wait where did you get the zero?

OpenStudy (anonymous):

ok yah haha

OpenStudy (anonymous):

thank you for helping me :)

myininaya (myininaya):

is that all you needed?

OpenStudy (anonymous):

after that would what would i do with my cscx (cscx-sinx)?

myininaya (myininaya):

I got 1/sinx-sinx+1=(1-[sinx]^2+sinx)/sinx=([cosx]^2+sinx)/sinx=[cotx]^2+1=(cscx)^2

myininaya (myininaya):

oops

OpenStudy (anonymous):

cool cool thank you so much is there a way to reward you or something by making me your fan or something?.... what?

myininaya (myininaya):

fan but I made a mistake

OpenStudy (anonymous):

how do i make fan.... and let me look at it ...

myininaya (myininaya):

do you see become a fan by my name?

OpenStudy (anonymous):

nope nothing is by your name

myininaya (myininaya):

no worries lets just find my mistake

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what did you do with the cscx tht was factored out?

myininaya (myininaya):

1/sinx-sinx+1 (here I found a common denominator) 1/sinx-(sinx)^2/sinx+sinx/sinx

myininaya (myininaya):

so we have (1-(sinx)^2+sinx)/sinx

OpenStudy (anonymous):

ok cool ..... is that all over sinx or just the sin^2 x +sinx?

myininaya (myininaya):

so we have ([cosx]^2+sinx)/sinx=[cosx]^2/sinx+sinx/sinx=(cosx)^2/sinx+1 BUT (cotx)^2+1=(cscx)^2. IS something missing from your question?

OpenStudy (anonymous):

my problem is cscx(cscx-sinx) + (sinx-cosx/sinx) +cotx = csc^2x

OpenStudy (anonymous):

so i think thats right ...

myininaya (myininaya):

omg im stupid i got it let me attach it

OpenStudy (anonymous):

it looks right except idk where you got the cos lol

myininaya (myininaya):

i dont know either

myininaya (myininaya):

and we could had done -1+1

myininaya (myininaya):

I made it way more difficult than it should had been sorry

OpenStudy (anonymous):

ok cool thanks i found the become a fan button

OpenStudy (anonymous):

and hey it worked out didnt it :P thanks a ton :)

myininaya (myininaya):

i will be ur fan too

OpenStudy (anonymous):

what do you get for having fans?

myininaya (myininaya):

a new car of course

OpenStudy (anonymous):

? a new car? haha what

myininaya (myininaya):

you get nothing but ego boosts i guess

OpenStudy (anonymous):

oh ok haha

OpenStudy (anonymous):

i guess it also shows how smart you are :)